OpenStudy (anonymous):

Solve x2 + 4x – 12 = 0 by completing the square. Show your work for full credit.

OpenStudy (amistre64):

it would be nice to add some useful form of zero to this thing to change the form, and not the value

OpenStudy (amistre64):

x2 + 4x + (0)– 12 = 0 x2 + 4x + (n^2 - n^2)– 12 = 0 (x2 + 4x + n^2) - n^2– 12 = 0 (x2 + 4x + n^2) = n^2 + 12 the issue being, what is a good value for n such that the left side is a perfect (or complete) square?

OpenStudy (amistre64):

nah, same generic variable would have been fine :) a perfect square is of the form: (x+n)^2 (x+n)^2 = (x+n) (x+n) = x^2 +2n x + n^2 there, now we have something to compare with x^2 + 4 x + n^2 x^2 + 2n x + n^2 so, what is the value of n if: 2n = 4?